OpenStudy (anonymous):

Which ordered pair in the form (a, b) is the solution of this equation? 4a - 5b = 24 A. (-5, -2) B. (6, 0) C. (-2, -5) D. (0, 6)

OpenStudy (anonymous):

4(-5)-5(-2) = -20+10 =10 4(6)-5(0) =24-0 =24 4(-2)-5(-5) =-8+25 =17 4(0)-5(6) =0-30 =-30

OpenStudy (anonymous):

i work out each options to see the result ..the one that give 24 at the end is ur option

OpenStudy (anonymous):

so (6, 0)

OpenStudy (anonymous):

yet

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

welcome